# RAVEN: Realtime Accessibility in Virtual ENvironments for Blind and Low-Vision People

### v1

@@ Abstract.
As virtual 3D environments become more prevalent, equitable access is essential for blind and low-vision (BLV) users, who face challenges with spatial awareness, navigation, and interaction. Prior work has explored supplementing visual information with auditory or haptic modalities, but these methods are static and offer limited support for dynamic, in-context adaptation. Recent advances in generative AI allow users to query and modify 3D scenes via natural language, introducing a paradigm that offers greater flexibility and control for accessibility. We present RAVEN, a system that enables BLV users to issue queries and modification prompts to improve the runtime accessibility of 3D virtual scenes. We evaluated RAVEN with eight BLV participants, generating empirical insights into how conversational programming can support personalized accessibility in 3D environments. Our work highlights both the promise of natural language interaction—intuitive, flexible, and empowering—and the challenges of ensuring reliability, transparency, and trust in generative AI–driven accessibility systems.
Accessibility, 3D, blind and low-vision
$c$s: Human-centered computing Accessibility systems and tools
Figure 1. System workflow of RAVEN, an interactive tool that enables BLV users to adapt 3D scenes through conversational natural language. The figure illustrates an example interaction: (A) A low-vision user types, “Make the sign with the text much bigger.” (B) The system combines semantic scene data with a runtime code-generation LLM to translate this request into an accessibility-enhancing modification. (C) The generated code is compiled and executed in real time, while the system also provides a spoken response confirming the change.:
@@ 1. Introduction
Virtual 3D environments have become pervasive, supporting a diverse range of applications such as interactive gaming, social networking, and education. These immersive spaces enable rich interactions, offering users a sense of presence and spatial exploration previously inaccessible in traditional media (Lim et al., 2006; Bowman et al., 2003). However, with their increased adoption arises the imperative of ensuring inclusive and equitable access, particularly for blind and low-vision (BLV) users who encounter challenges in spatial understanding, navigation, and object interaction in these environments (White et al., 2008).
To address these challenges, prior work proposed tools that modify or supplement visual information using alternative modalities, such as audio descriptions (Gonzalez-Mora et al., 2006), haptic feedback (Jansson et al., 1999; Zhao et al., 2018a), and enhanced visual effects tailored for low vision (Zhao et al., 2019b). Systems like SceneWeaver (Balasubramanian et al., 2023) offer users greater agency over when and how to consume scene descriptions. Commercial platforms have also adopted accessibility features—such as high-contrast display modes and spatial audio—to accommodate BLV players in mainstream video games (Bayliss, 2022; Playstation, 2024). These methods allow BLV users to utilize pre-defined tools to enhance the accessibility of a virtual 3D environment.
@@
We evaluated RAVEN with eight BLV participants in a user study including three interactive scenarios: a guided tutorial, structured tasks, and free exploration. Participants engaged in conversational interactions to tailor scenes according to their accessibility preferences, offering quantitative feedback on usability and qualitative reflections on their experiences. Findings reveal critical insights into the potential and limitations of generative AI-supported interactions, highlighting both opportunities for enhanced accessibility and challenges with trust and reliability.
In summary, our work contributes: (1) the design and implementation of RAVEN, a GenAI-powered system for real-time, natural language-driven querying and modification of 3D environments for accessibility, and (2) empirical insights from a study with eight BLV users, highlighting interaction strategies, usability, and challenges that inform future LLM-based accessibility tools. Our work demonstrates the importance of flexible, intuitive, and contextually adaptive accessibility in virtual environments.
@@ 2.1. BLV Accessibility in Virtual 3D Environments
Early efforts created bespoke virtual environments for BLV users (SÁNCHEZ and LUMBRERAS, 1999; Gonzalez-Mora et al., 2006; Picinali et al., 2014; Andrade et al., 2018), for activities including rehabilitation and orientation training programs (Seki and Sato, 2011; Tzovaras et al., 2009) and specific experiences such as boxing (Furtado et al., 2025), racing (Gluck et al., 2021), and table tennis (Kamath et al., 2024). While these systems demonstrate feasibility and value, their design assumptions often limit generalizability beyond the particular contexts for which they were built.
@@ 2.2. Generative AI for Visual Accessibility
GenAI also enables visual content creation, spanning images (Zhao et al., 2024; Bordes et al., 2024) and 3D scenes (Hu et al., 2024; Zhang et al., 2024), opening avenues for creative agency among BLV users (Bennett et al., 2024). Yet generated content is often inaccessible to its creators, complicating evaluation, verification, and iterative refinement (Huh et al., 2023). To address this in 2D settings, GenAssist combines vision-language and language models to pose verification and style questions, and synthesize question-guided descriptions that help users assess prompt alignment and compare candidates (Huh et al., 2023). AltCanvas introduces a spatially structured, tile-based interface to improve layout comprehension while adding, editing, and moving visual components (Lee et al., 2024). While effective in 2D, analogous accessibility support for virtual 3D generative workflows remains underexplored.
@@ 3. Iterative System Design
Figure 2. Examples of system usage. The center image shows the original game scene, with six surrounding panels illustrating accessibility-driven modifications: (A) change object color, (B) reposition the player, (C) enlarge a text object, (D) increase brightness, (E) amplify audio volume, and (F) adjust audio pitch. Bubbles and icons in E and F visualize auditory changes. These are only a few of the many types of modifications RAVEN can support to enable flexible, user-driven accessibility.:
RAVEN supports real-time queries and modifications in virtual 3D environments to improve accessibility for BLV users. To ground its features in user needs, we first developed an initial prototype and conducted a pilot study with three BLV participants. Insights from this study informed iterative refinements, resulting in the final system presented here. In this section, we introduce an illustrative use scenario (section 3.1), describe the pilot study and resulting design modifications (section 3.2), and detail the final system architecture (section 3.3).
@@ 3.3. The RAVEN System
The final RAVEN architecture (Figure 1-B) requires only minimal developer annotation during scene creation (section 3.3.1) and integrates several components to support natural language interaction. At runtime, the system self-voices user input (section 3.3.2), retrieves an accessibility-augmented semantic scene graph (SSG) from the environment (section 3.3.3), and constructs prompts that combine user queries with accessibility and error-prevention instructions (section 3.3.4). These prompts are executed through the GROMIT runtime generation system (Jennings et al., 2024), which generates Unity code to implement requested modifications and returns a textual response. This response is announced to the user through text-to-speech, completing the interaction loop. The full system implementation is open sourced: [link omitted for review].
@@ 3.3.1. Scene Construction
To prepare a scene, developers create a 3D environment in Unity following standard workflows, then identify a subset of important objects (e.g., , game items on a table). For each selected object, a Unity script is attached that specifies metadata, including the object’s name, a textual description of its visual and auditory properties, and whether it is physical (e.g., , a table) or non-physical (e.g., , an ambient sound source). Developers also include lightweight scripts that connect the scene to the LLM agent. This metadata is the only additional input required, keeping developer burden minimal.
@@ 3.3.3. Dynamic Information Retriever
The Dynamic Information Retriever produces an up-to-date semantic scene graph (SSG) at each user query. Building on the structure introduced by Jennings et al. (Jennings et al., 2024), we extend the SSG with accessibility metadata. For each relevant object, the SSG stores its name, developer-provided description (visual, auditory, and functional properties), attached scripts, position, scale, and hierarchical relationships. To support accessibility queries and modifications, we further augment this data with:
@@ 3.3.4. Prompt Constructor
1. Accessibility Support Instructions. Although LLMs often possess accessibility knowledge, they may struggle to apply it effectively in specific contexts (Glazko et al., 2023). To address this, we embedded a set of rules tailored to the Unity3D environment, grounded in prior work on BLV accessibility needs in 3D environments (White et al., 2008; Nair et al., 2021; Guerreiro et al., 2023). These rules guide the LLM in providing scene descriptions and applying modifications related to navigation, object identification, sound, color and lighting, and text size.
2. Error Prevention Instructions. To mitigate hallucinations and improve transparency, we designed instructions that enable the system to handle ambiguity and errors conversationally. If a request is incomplete or vague, the LLM asks clarifying questions (e.g., “It seems like your request is not clear. Could you please provide more details or clarify what you would like to achieve?”). If a user reports an error (e.g., , “it’s not working”), the system is instructed to acknowledge the limitation and suggest alternative approaches, for instance: “I apologize for the error. Let’s try a different approach to stop the Apple from moving.” These strategies allow the system to engage in follow-up dialogue, helping users refine prompts or explore different accessibility adjustments. Finally, the LLM is primed (Sun et al., 2025) with a list of out-of-scope tasks identified during the pilot study (e.g., adding magnifiers). This ensures the system informs users when a request is unsupported rather than attempting a potentially erroneous modification.
@@ 3.3.5. Runtime Modification Generation and Compilation
Once the prompt is constructed, it is sent to GROMIT (NicholasJJ, 2024; Jennings et al., 2024), an open-source runtime behavior generation system for Unity. GROMIT processes the prompt by generating Unity code that implements the requested modification, attaches the compiled script to the relevant object, and returns a textual response. In RAVEN, we upgraded GROMIT to use GPT-4o for improved language and code generation. This enables the system to dynamically update scenes while providing users with immediate, self-voiced feedback.
@@ 4.1. Accessibility Categories
Guided by prior research on accessibility for BLV users in virtual environments (Szpiro et al., 2016; Wurm et al., 1993; Rigden, 1999; Uysa and Düger, 2012; Rubin et al., 2006; Zhao et al., 2019b), we defined six categories of queries and modifications to structure the evaluation. Four address the visual domain and two target the auditory dimension.
@@ 4.2. Scenes
We designed three scenes to scaffold participants’ exploration of the six accessibility categories, with increasing complexity and freedom of interaction (Figure 3).
@@ 4.2.1. Scene 1: Guided Tutorial
The first scene introduced participants to the system in a simple, game-like room containing two geometric objects with distinct colors, two speakers playing dialogue, a torch with continuous sound and animation, a text object, and a table (fig. 3(a)). The researcher demonstrated each category using a three-step prompt pattern (ask, modify, verify), after which participants practiced similar prompts themselves. This structure, inspired by prior work on generative AI accessibility tools (Glazko et al., 2023; Adnin and Das, 2024; Huh et al., 2023), ensured that participants could explore both querying and modification. Full list of demo prompts used are included in section A.2.
@@ 4.2.2. Scene 2: Task-Driven Exploration
The second scene, a cat park with natural elements, three cats producing distinct meows, a bench, a streetlamp, and a garden hut (fig. 3(b)), was paired with six predefined tasks (table 1). The tasks were designed to align with the six accessibility categories. After completing a task, participants reflected on their experience, including whether the interaction was intuitive, useful, or confusing. This setup allowed us to observe how participants applied categories in a goal-oriented context.
@@ 4.2.3. Scene 3: Open-Ended Exploration
The final scene encouraged open-ended use of the system. Participants explored a spaceship-themed room containing sixteen objects and three sound sources (fig. 3(c)). They were given ten minutes to query, modify, and adapt the environment according to their own needs, without predefined tasks. This scenario revealed how participants appropriated the system in less structured contexts.
(a) Scene 1:
(b) Scene 2:
(c) Scene 3:

### v2

@@ Abstract.
As virtual 3D environments become more prevalent, equitable access is essential for blind and low-vision (BLV) users, who face challenges with spatial awareness, navigation, and interaction. Prior work has explored supplementing visual information with auditory or haptic modalities, but these methods are static and offer limited support for dynamic, in-context adaptation. Recent advances in generative AI allow users to query and modify 3D scenes via natural language, introducing a paradigm that offers greater flexibility and control for accessibility. We present RAVEN, a system that enables BLV users to issue queries and modification prompts to improve the runtime accessibility of 3D virtual scenes. We evaluated RAVEN with eight BLV people and six Unity developers, generating empirical insights into how conversational programming can support personalized accessibility in 3D environments. Our work highlights both the promise of natural language interaction—intuitive, flexible, and empowering—and the challenges of ensuring reliability, transparency, and trust in generative AI–driven accessibility systems.
Accessibility, 3D, blind and low-vision, generative AI
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791616isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Accessibility systems and tools
System workflow showing a BLV user modifying a 3D scene. On the left, a low-vision user types the request ”Make the sign with the text much bigger” while viewing a game scene. In the center, the system processes the scene objects and user prompt. On the right, the updated scene shows a larger sign, and a text-to-speech message confirms the change.
Figure 1. System workflow of RAVEN, an interactive tool that enables BLV users to adapt 3D scenes through conversational natural language. The figure illustrates an example interaction: (A) A low-vision user types, ‘Make the sign with the text much bigger.” (B) The system combines semantic scene data with a runtime code-generation LLM to translate this request into an accessibility-enhancing modification. (C) The generated code is compiled and executed in real time, while the system also provides a spoken response confirming the change.
@@ 1. Introduction
Virtual 3D environments have become pervasive, supporting a diverse range of applications such as interactive gaming, social networking, and education. These immersive spaces enable rich interactions, offering users a sense of presence and spatial exploration previously inaccessible in traditional media (Lim et al., 2006; Bowman et al., 2003). However, with their increased adoption arises the imperative of ensuring inclusive and equitable access, particularly for blind and low-vision (BLV) users who encounter challenges in spatial understanding, navigation, and object interaction in these environments (White et al., 2008). s To address these challenges, prior work proposed tools that modify or supplement visual information using alternative modalities, such as audio descriptions (Gonzalez-Mora et al., 2006), haptic feedback (Jansson et al., 1999; Zhao et al., 2018a), and enhanced visual effects tailored for low vision (Zhao et al., 2019b). Systems like SceneWeaver (Balasubramanian et al., 2023) offer users greater agency over when and how to consume scene descriptions. Commercial platforms have also adopted accessibility features—such as high-contrast display modes and spatial audio—to accommodate BLV players in mainstream video games (Bayliss, 2022; Playstation, 2024). These methods allow BLV users to utilize pre-defined tools to enhance the accessibility of a virtual 3D environment.
@@
We evaluated RAVEN with eight BLV participants in a user study including three interactive scenarios: a guided tutorial, structured tasks, and free exploration. The scenarios and study tasks were designed around six accessibility categories drawn from prior work. Participants engaged in conversational interactions to tailor scenes according to their accessibility preferences, offering quantitative feedback on usability and qualitative reflections on their experiences. Findings reveal critical insights into the potential and limitations of generative AI-supported interactions, highlighting both opportunities for enhanced accessibility and challenges with trust and reliability. We also conducted a preliminary study with six Unity developers to evaluate the system’s usability from a developer perspective, yielding insights into its learnability, overall usability, and opportunities for improvement.
In summary, our work contributes: (1) the design and implementation of RAVEN, a GenAI-powered system for real-time, natural language-driven querying and modification of 3D environments for accessibility, and (2) empirical insights from a study with eight BLV users and a preliminary study with six Unity developers, highlighting interaction strategies, usability, and challenges that inform future LLM-based accessibility tools. Our work demonstrates the importance of flexible, intuitive, and contextually adaptive accessibility in virtual environments.
@@ 2.1. BLV Accessibility in Virtual 3D Environments
For the context of this work, virtual 3D environments are computer-generated three-dimensional spaces that may be displayed on PCs, mobile devices, or immersive VR headsets. Early efforts created bespoke virtual environments for BLV users (SÁNCHEZ and LUMBRERAS, 1999; Gonzalez-Mora et al., 2006; Picinali et al., 2014; Andrade et al., 2018), for activities including rehabilitation and orientation training programs (Seki and Sato, 2011; Tzovaras et al., 2009) and specific experiences such as boxing (Furtado et al., 2025), racing (Gluck et al., 2021), and table tennis (Kamath et al., 2024). While these systems demonstrate feasibility and value, their design assumptions often limit generalizability beyond the particular contexts for which they were built.
@@
Despite these advances, most approaches remain developer-defined and static: substitutions (e.g., auditory cues) and modifications (e.g., color mappings) are predetermined and may not track the diverse, evolving needs of BLV users (Creed et al., 2024). They can also impose learning burdens through mode switches, keybindings, or complex settings, which is especially challenging for newcomers (Kreimeier and Götzelmann, 2020; Maidenbaum et al., 2016).
We also build on a prior non-archival ASSETS demo that presented an early prototype based on our work  (Cao et al., 2025). That demo introduced the idea of using LLM-driven runtime modifications in Unity and reported preliminary results, but it did not include the full analysis of the BLV user study or the developer study reported here. In this paper, we substantially extend that prototype by (1) fully specifying the system architecture and prompt-engineering framework (section 3.3.1-section 3.3.6), (2) deriving and documenting accessibility rules and categories from BLV literature, and (3) contributing two new empirical evaluations: a three-scene study with eight BLV participants and a preliminary usability study with six Unity developers. We therefore treat the demo as an early proof-of-concept and this paper as the first archival presentation of the complete system and its evaluation.
To our knowledge, RAVEN is the first system to enable BLV users to query and modify virtual 3D scenes at runtime via natural language, shifting control from developer-defined presets to user-directed, in-situ adaptation.
@@ 2.2. Generative AI for Visual Accessibility
GenAI also enables visual content creation, spanning images (Zhao et al., 2024; Bordes et al., 2024) and 3D scenes (Hu et al., 2024; Zhang et al., 2024), opening avenues for creative agency among BLV users (Bennett et al., 2024). Yet generated content is often inaccessible to its creators, complicating evaluation, verification, and iterative refinement (Huh et al., 2023). To address this in 2D settings, GenAssist combines vision-language and language models to pose verification and style questions, and synthesize question-guided descriptions that help users assess prompt alignment and compare candidates (Huh et al., 2023). AltCanvas introduces a spatially structured, tile-based interface to improve layout comprehension while adding, editing, and moving visual components (Lee et al., 2024). While effective in 2D, analogous accessibility support for virtual 3D generative workflows remains underexplored, where BLV users face additional demands for spatial and embodied understanding.
@@ 3. Iterative System Design
A collage of images showing video game scenes. One image is in the center, pointing arrows towards six images that surround it. The texts on the arrows are commands: Make the car bright yellow; Move me next to the policeman; Make the car plate bigger; Make the sunlight brighter; Make the pedestrian louder; Increase policeman’s pitch. The six surrounding images show the corresponding changes.
Figure 2. Examples of system usage. The center image shows the original game scene, with six surrounding panels illustrating accessibility-driven modifications: (A) change object color, (B) reposition the player, (C) enlarge a text object, (D) increase brightness, (E) amplify audio volume, and (F) adjust audio pitch. Bubbles and icons in E and F visualize auditory changes. These are only a few of the many types of modifications RAVEN can support to enable flexible, user-driven accessibility.
RAVEN supports real-time queries and modifications in virtual 3D environments to improve accessibility for BLV users. The system is designed to be open, where users can use any natural language prompt. However, to ground its utility in user needs, we first developed an initial prototype and conducted a pilot study with three BLV participants. Insights from this study informed iterative refinements, resulting in the final system presented here. In this section, we introduce an illustrative use scenario (section 3.1), describe the pilot study and resulting design modifications (section 3.2), and detail the final system architecture (section 3.3).
@@ 3.3. The RAVEN System
The final RAVEN architecture (fig. 1-B) requires only minimal developer annotation during scene creation (section 3.3.1) and integrates several components to support natural language interaction. At runtime, the system self-voices user input (section 3.3.2), retrieves an accessibility-augmented semantic scene graph (SSG) from the environment (section 3.3.3), and constructs prompts that combine user queries with accessibility and error-prevention instructions (section 3.3.4). These prompts are executed through the GROMIT runtime generation system (Jennings et al., 2024), which generates Unity code to implement requested modifications and returns a textual response. This response is announced to the user through text-to-speech, completing the interaction loop. The full system implementation is open-sourced and can be accessed here: https://github.com/SoundabilityLab/RAVEN.
@@ 3.3.1. Scene Construction
To prepare a scene, developers create a 3D environment in Unity following standard workflows, then identify a subset of important objects (e.g., game items on a table). For each selected object, developers attach a lightweight Unity script and indicate whether it is physical (e.g., a table) or non-physical (e.g., an ambient sound source), and whether it is the player. Optional developer-provided descriptions of visual or auditory properties may also be included. Developers also include lightweight scripts that connect the scene to the LLM agent. This metadata is the only additional input required, keeping developer burden minimal.
@@ 3.3.3. Dynamic Information Retriever
The Dynamic Information Retriever produces an up-to-date semantic scene graph (SSG) each time a prompt is issued, ensuring that the LLM receives current environmental context. Building on the structure introduced by Jennings et al. (Jennings et al., 2024), we extend the SSG with accessibility metadata. For each relevant object, the SSG stores its name, developer-provided description (visual, auditory, and functional properties), attached scripts, position, scale, and hierarchical relationships. To support accessibility queries and modifications, the following metadata is recomputed at every prompt:
@@ 3.3.4. Prompt Constructor
1. Accessibility Support Instructions. Although LLMs may possess accessibility knowledge, they may struggle to apply it effectively in specific contexts (Glazko et al., 2023). We therefore synthesized accessibility needs in virtual 3D environments from prior BLV accessibility research (White et al., 2008; Nair et al., 2021; Guerreiro et al., 2023; Zhao et al., 2019b). These needs span color/contrast adjustments, brightness changes, spatial understanding, text enlargement, scene description, and audio manipulation. We operationalized these needs into a set of explicit rules (e.g., how to adjust Unity C# variables, how to simplify materials before recoloring, and how to convert Unity terminology into user-facing language) that guide the LLM’s descriptions and modifications.
2. Error Prevention Instructions. To mitigate hallucinations and improve transparency, we designed instructions that enable the system to handle ambiguity and errors conversationally. If a request is incomplete or vague, the LLM asks clarifying questions (e.g., “It seems like your request is not clear. Could you please provide more details or clarify what you would like to achieve?”). If a user reports an error (e.g., “it’s not working”), the system is instructed to acknowledge the limitation and suggest alternative approaches. Finally, the LLM is primed (Sun et al., 2025) with a list of out-of-scope tasks identified during the pilot study (e.g., adding magnifiers), ensuring that unsupported requests are clearly surfaced rather than producing erroneous code.
@@ 3.3.5. Runtime Modification Generation and Compilation
Once the prompt is constructed, it is sent to GROMIT (NicholasJJ, 2024; Jennings et al., 2024), an open-source runtime behavior generation system for Unity. GROMIT processes the prompt by generating Unity code that implements the requested modification, attaches the compiled script to the relevant object, and returns a textual response. This enables the system to dynamically update scenes while providing users with immediate, self-voiced feedback. We used GPT-4o for language and code generation, as it was the most advanced model that supported feasible real-time integration during system development.
@@ 4.1. Accessibility Categories
To scaffold our study design, we synthesized six accessibility categories from prior work across game accessibility toolkits and empirical studies of BLV user needs (Szpiro et al., 2016; Wurm et al., 1993; Rigden, 1999; Uysa and Düger, 2012; Rubin et al., 2006; Zhao et al., 2019b). These categories were used solely to structure our evaluation tasks and analysis; they do not constrain RAVEN’s capabilities nor restrict what users may request during interaction. Four categories address the visual domain and two target the auditory dimension.
@@ 4.2. Scenes
We designed three scenes to scaffold participants’ exploration of the six accessibility categories, with increasing complexity and freedom of interaction (fig. 3).
@@ 4.2.1. Scene 1: Guided Tutorial
The first scene introduced participants to the system in a simple, game-like room containing two geometric objects with distinct colors, two speakers playing dialogue, a torch with continuous sound and animation, a text object, and a table (fig. 3a). The researcher demonstrated each category using a three-step prompt pattern (ask, modify, verify), after which participants practiced similar prompts themselves. This structure, inspired by prior work on generative AI accessibility tools (Glazko et al., 2023; Adnin and Das, 2024; Huh et al., 2023), ensured that participants could explore both querying and modification. Full list of demo prompts used are included in section A.2.
@@ 4.2.2. Scene 2: Task-Driven Exploration
The second scene, a cat park with natural elements, three cats producing distinct meows, a bench, a streetlamp, and a garden hut (fig. 3b), was paired with six predefined tasks (table 1). The tasks were designed to align with the six accessibility categories. After completing a task, participants reflected on their experience, including whether the interaction was intuitive, useful, or confusing. This setup allowed us to observe how participants applied categories in a goal-oriented context.
@@ 4.2.3. Scene 3: Open-Ended Exploration
The final scene encouraged open-ended use of the system. Participants explored a spaceship-themed room containing sixteen objects and three sound sources (fig. 3c). They were given ten minutes to query, modify, and adapt the environment according to their own needs, without predefined tasks. This scenario revealed how participants appropriated the system in less structured contexts.
An image of a video game scene. There is a white table with a green sphere, a red cube and a burning torch on top of it. Some text is hovering over the table.
(a) Scene 1
An image of a video game scene in a park. There is a road extending from left to right into the distance. A bench under a street lamp is next to the road. Several cats are on the road. There’s a garden hut in the distance.
(b) Scene 2
An image of an indoor video game scene. There are two tables in the room. The table on the right has colorful vases and small items on top. There’s a piece of white text on the wall above this table. The table on the left is surrounded by colorful chairs and has small items on top.
(c) Scene 3